OpenStudy (anonymous):

At time t hours after taking the cough suppressant hydrocodone bitartrate, the amount, A, in mg, remaining in the body is given by A = 14(0.8)t. What percent of the drug leaves the body each hour? How much of the drug is left in the body 2 hours after the dose is administered? How long is it until only 1 mg of the drug remains in the body?

mathslover (mathslover):

first of all allow me to know that what you tried?

OpenStudy (anonymous):

.8+1=1.8 so .018%

OpenStudy (anonymous):

but thats wrong

OpenStudy (anonymous):

Do you actually mean \(A=14\cdot 0.8\cdot t\), or did you mean \(A=14t^.8\)?

OpenStudy (anonymous):

A=14(0.8)^t

OpenStudy (anonymous):

\[14(0.8)^T\]

mathslover (mathslover):

@asnaseer may help u here.. sorry

OpenStudy (asnaseer):

the amount of drug in the body after t hours is given by:\[A=14(0.8)^t\]so, when t=0 (i.e. at the start), the amount in the body is given by:\[A=14(0.8)^0=14mg\]after each hour, the amount left in the body is multiplied by 0.8, which represents 80%. so how much drug leaves the body each hour?

OpenStudy (asnaseer):

@Ldaniel I can only help if you respond to my questions.

OpenStudy (anonymous):

20%

OpenStudy (anonymous):

@asnaseer